1.find perimeter of a square with diagonal of 12 centimeters
2.find the diagonal of a square with a perimeter of 20 inches
3.find the diagonal of a square with a perimeter of 28 inches

Answer by pug912(10)

1.find perimeter of a square with diagonal of 12 centimeters
Let d be the lenght of every side of a square. then the perimeter = 4*d centimeters
Using Pythagoras Theorem, x^2+y^2 = z^2
where x=y=d and z = diagonal length = 12
d^2+d^2=12^2
2*d^2 = 144
d^2 = 144/2 = 72
so d = sqrt(72)
so perimeter = 4*sqrt(72)
2.find the diagonal of a square with a perimeter of 20 inches
since perimeter = 20, so each side of square = 20/4 = 5 inches
Let d be the lenght of diagonal.
Using Pythagoras Theorem, x^2+y^2 = z^2
where x=y=5 and z=d
5^2+5^2 = d^2
d^2 = 25+25 = 50
sp d = sqrt(50) inches = length of diagonal
3.find the diagonal of a square with a perimeter of 28 inches
since perimeter = 28, so each side of square = 28/4 = 7 inches
Let d be the lenght of diagonal.
Using Pythagoras Theorem, x^2+y^2 = z^2
where x=y=7 and z=d
7^2+7^2 = d^2
d^2 = 49+49 = sqrt(98)
sp d = sqrt(98) inches = length of diagonal